SUB-DOPPLER MEASUREMENTS ON THE ROTATIONAL TRANSITIONS OF CARBON MONOXIDE

G. WINNEWISSER, \underline{S. P. BELOV}, TH. KLAUS, and R. SCHIEDER, I. Physikalisches Institut, Universität zu Köln, Zülpicher Str. 77, D-50937 Köln, Germany.

The frequencies of the first six rotational transitions of the CO molecule have been measured to an accuracy of ±500 Hz in the frequency region up to 700 GHz. This high level of accuracy was achieved with the Cologne terahertz spectrometer operated in the sub-Doppler mode (1). To carry out precise CO Lamb-dip measurements, widely tunable BWO’s (OB-30, OB-32, and OB-80), phase-locked to a KVARZ synthesizer (78-118 GHz), were used. Least squares fits of the new CO data separately and together with our previous Doppler-limited measurements up to 1.3 THz (2) yielded rms values of 0.24 kHz and 0.26 kHz, respectively. The experimental setup and the results of a fit to the Lamb-dip data combined with the frequencies up to 4.3 THz (J = 38 ← 37) given by Evenson (3) will be presented. The analysis yielded a revised set of rotational constants for CO in its vibrational ground state, which is in good agreement with the previous values given by Varberg and Evenson (4).

(1) G. Winnewisser, S.P. Belov, Th. Klaus, and Š Urban, Z. Naturforsch. 51a, 200-206 (1996).

(2) S.P. Belov, F. Lewen, Th. Klaus, and G. Winnewisser, J. Mol. Spectrosc. 174, 606-612 (1995).

(3) K.M. Evenson, private communication, 52 Okazaki Conference, Okazaki, Japan, March 1995.

(4) T.D. Varberg and K.M. Evenson, Astrophys. J. 385, 763-765 (1992).